OpenStudy (anonymous):

really simple interest question!! Lily deposited $19 in a savings account that earns 1.3% simple interest. Which graph represents this scenario?

OpenStudy (anonymous):

@ganeshie8 can you explain really quick?

OpenStudy (anonymous):

I think it is the 3rd option. the first graph in the second picture

OpenStudy (nevermind_justschool):

Dang it, I wish I could help but I'm not allowed to go to unknown sites. Plus, I'm not very good at interest. Sorry! Good luck to ya!

OpenStudy (anonymous):

i got nothing, wish i could help but I'm not the greatest at math

OpenStudy (anonymous):

hahah well thanks for trying!

OpenStudy (k_lynn):

@mathmate @mathstudent55 could probably help

OpenStudy (anonymous):

okay, thanks!

OpenStudy (mathmate):

hint: the total amount (balance) after x years with a principal of $19 and with simple interest of 1.3% is given by: balance(x) = 19(1+0.013x) Calculate for two values of x (for example x=0, balance = 19(1+0)=$19 and choose the graph that fits both points.